OpenStudy (anonymous):

The equation of the normal line to the curve y = cube root (x^2 - 1) at the point where x = 3 is

OpenStudy (anonymous):

Take the derivative, find the slope, take the negative reciprocal to find the orthogonal slope, then plug in the point to find the y coordinate, then use point slope form y-y_0=m(x-x_0) to find your line.

OpenStudy (anonymous):

\[y=\sqrt[3]{x^2-1}\] \[y'=(x^2-1)^{-\frac{2}{3}}*2x\] The slope of the normal line is the negative reciprocal of this, so at (3,2)=(x,y) y'=24 Normal line slop = -1/24 Put this into point-slop form for a line: y-2=(-1/24)(x-3)

OpenStudy (anonymous):

I'm a bit confused after we took the derivative... where is the (3,2) coming from?

OpenStudy (anonymous):

Just plug 3 into the original function. It is the point we are finding the normal at.

OpenStudy (anonymous):

So what isn't making sense?

OpenStudy (anonymous):

the reciprocal part

OpenStudy (anonymous):

For any line: y=mx The line: y=-(1/m)x is perpendicular / normal /orthogonal where they intersect. Try playing with it a bit, sketching it.

OpenStudy (anonymous):

sorry, now I dont see how y' = 24

OpenStudy (anonymous):

Just calculate y'(x) where x=3

OpenStudy (anonymous):

I keep getting 3/2

OpenStudy (anonymous):

You are correct. I forgot to put a negative sign in my exponent above, and haven't checked my work. it is 3/2.